Fermi National Laboratory


Accelerator Update

Wednesday June 25
The day shift began with Tevatron (TeV) experts investigating the cause of the sector BA quenches, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

Booster dampers tripped off at 9:56 AM, which also tripped off MiniBooNE. An expert investigated and had the dampers back on by 11:14 AM.

MiniBooNE tripped off at 1:54 PM due to a quadrupole power supply (Q874) breaker problem. The Duty Electrician (DE) reset the breaker.

Operations put the TeV through a dry squeeze at 2:53 PM and then put the TeV into shot setup at 3:13PM.

At 5:13 PM, Operations established store 2723 with an initial luminosity of 30.61E30. Pbar resumed stacking.

A Booster dipole power supply (B:V803) tripped off on high temperature at 6:18 PM. Operators put a fan on the supply and had it back in operation by 6:47 PM.

MiniBooNE's Q874 tripped off at 6:25 PM. The DE reset the breaker and Operations resumed sending beam to MiniBooNE at 7:03 PM. It then tripped off again at 8 PM and then three more times over the next three hours.

Thursday June 26
The midnight shift began with Operations monitoring store 2723, with Pbar stacking, and with MiniBooNE taking beam.

Q874 tripped again at 12:09 AM. The DE reset the breaker at 12:35 AM.

At 7:40 AM, Operations turned off MiniBooNE beam so a tech could replace the Q874 breaker. Beam returned to MiniBooNE at 8:15 AM.

A Switchyard expert began a study period at 8:53 AM.

Operations terminated store 2723 at 10:35 AM. Shot setup began at 11:30 AM.

Operations established store 2725 at 1:20 PM. Its initial luminosity was 33.85E30. Pbar resumed stacking.

Operations began conducting some NuMI studies at 3:42 PM.

Another Switchyard expert began a study period that ran from 7 PM until a little after midnight.

Friday June 27
The midnight shift began with Operations monitoring store 2725, with Pbar stacking, and with MiniBooNE taking beam.

A Main Injector expert conducted a study from 12:34 AM to 3:24 AM.

Two Booster RF stations (BRF5 & 6) tripped off at 4:52 AM due to a vacuum burst. The same trip occurred three more times over the next fifteen minutes.

Today's Plan
The plan for today and this weekend is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/27/2003   email Fermilab